# Random Chowla’s Conjecture for Rademacher Multiplicative Functions

### v1

@@ Abstract.
We study the distribution of partial sums of Rademacher random multiplicative functions $(f(n))_{n}$ evaluated at polynomial arguments. We show that for a polynomial $P\in\operatorname{\mathbb{Z}}[x]$ that is a product of distinct linear factors or an irreducible quadratic satisfying a natural condition, there exists a constant $\kappa_{P}>0$ such that
@@ 1.1. Random Multiplicative Functions
Note that $f(n)$ is a random model for $\mu(n),$ as $\mu(n)$ is simply a specific realization of $f(n)$. Using the theory of Dirichlet series, Wintner [Win44] showed that $\sum_{n}f(n)n^{-s}$ is almost always convergent for $\Re(s)>1/2.$ He further showed that, for all $\varepsilon>0$, both $\sum_{n\leq x}f(n)=O_{\varepsilon}(x^{1/2+\varepsilon})$ and $\sum_{n\leq x}f(n)\neq O_{\varepsilon}(x^{1/2-\epsilon})$ hold almost always, leading the egregious statement that “the Riemann Hypothesis is almost surely true”.
@@ 1.2. The Random Chowla Conjecture
Let $f$ be a Rademacher RMF and let $P\in\operatorname{\mathbb{Z}}[x]$ be a product of distinct linear factors over $\operatorname{\mathbb{Z}}$ or irreducible of degree $2$. Suppose that there is no prime $p$ for which $p^{2}\mid P(n)$ for all $n\in\operatorname{\mathbb{N}}$. Then, there exists a constant $\kappa_{P}>0$ such that
@@
respectively. The former is still “trivial”, as a product of two squarefree numbers is a perfect square if and only if the two numbers are equal; and so, the second moment amounts to counting the number of $n\leq N$ for which $P(n)$ is squarefree. This is a notoriously difficult problem for general polynomials, which we briefly discuss at the end of Section 2. The main technical result in this paper is proving an asymptotic for (1.3) (Proposition 3.1). The key new ingredient to our proof is a bootstrapping argument, where it is essential that our polynomials are products of linear factors or simply irreducible quadratics.
@@ 1.4. Completely Multiplcative RMFs and Quadratic Twists
which counts the number of integral points up to a given height on a family of twists of a hyperelliptic curve with affine model $C_{d}:dy^{2}=P(x)$. In the case where $P$ has degree $2$, we are able to solve this counting problem using work of Hooley on Pell equations [Hoo84]: Theorem 1 of [Hoo84] tells us that almost all twists will have at most $1$ integral point of low height, so that we can replace the second moment for these $d$ by the first moment, which exactly counts diagonal solutions $n_{1}=n_{2}$. For the remaining twists, trivial bounds suffice. For higher degree polynomials, analogues of Hooley’ results exist, conditionally on the $abc$-conjecture and for suitable $P$, by work of Granville [Gra07, Theorem 4 (i)]; as such, we are able to solve special cases of the second moment problem for completely multiplicative Rademacher RMFs. The fourth moment is substantially more complicated. For starters, one would like to know that almost all twists of the surface $y^{2}=P(x_{1})P(x_{2})$ have at most one integral point. Nevertheless, we still expect the behaviour to be Gaussian and this is a subject of ongoing investigation.
@@ 2.1. Martingales and McLeish’s CLT
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $(\mathcal{F}_{n})_{n}$ be a filtration on $(\Omega,\mathcal{F},\mathbb{P})$ (i.e., a sequence of sub-$\sigma$-algebras satisfying $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\dots\subseteq\mathcal{F}$). A sequence of random variables $(X_{n})_{n}$ is called a martingale difference sequence on $(\Omega,\mathcal{F},(\mathcal{F}_{n})_{n},\mathbb{P})$ iff
@@ 2.2. Squarefree Values and Divisor Bounds
Let $Q\in\operatorname{\mathbb{Z}}[x]$, $y,N\in\operatorname{\mathbb{N}}$, and $\omega(y):=\#\{p:p\mid y\}$. If $y$ is both squarefree and coprime to the leading coefficient of $q$, then:
@@
Given $F(x,y),$ consider its prime factorization $f_{1}(x,y)\cdots f_{k}(x,y)$ in $\operatorname{\mathbb{C}}[x,y]$ and note that $F(x,y)=0$ if and only if $f_{i}(x,y)=0$ for some $i$; in particular, the number of integral points on $F(x,y)=0$ is the sum of the number of integral points on each $f_{i}(x,y)$. For $f_{i}$ with $\deg f_{i}\geq 3$, there are at most $\ll_{\varepsilon}N^{1/\deg f_{i}+\varepsilon}\ll N^{1/3+\varepsilon}$ integral points in the desired box, by Lemma 2.3. For $f_{i}$ of degree $2$, we may assume, w.l.o.g., that $f_{i}(x,y)\in\operatorname{\mathbb{Q}}[x,y]$: if not, we note that $f_{i}$ comes with a conjugate pair, say $f_{j}$, as $F(x,y)\in\operatorname{\mathbb{Z}}[x,y];$ in particular, the number of rational points on $f_{i}(x,y)=0$ is equal to the number of rational points on the intersection of the curves $f_{i}(x,y)=0=f_{j}(x,y)$, which is at most $4$ by Bezout’s Theorem on intersections of plane curves [Béz79]. In the case where $F(x,y)$ has quadratic factors in $\operatorname{\mathbb{Q}}[x,y]$, then the coefficients of this quadratic factor are $\ll N^{O_{\deg P}(1)}$ whenever $a,b\ll N^{O_{\deg P}(1)}$, so that there are at most $\ll_{\varepsilon}N^{\varepsilon}$ integral points across the quadratic irreducible factors (by Lemma 2.4). Showing the discriminants are non-square and that the coefficients of the quadratic factors are bounded by $N^{O_{\deg P}(1)}$ is a relatively simple task: one can do this directly by noticing that the LHS of $F(x,y)=f_{1}(x,y)\cdots f_{k}(x,y)$ is at most $N^{O_{\deg P}(1)}$ for $a,b\ll N^{O_{\deg P}(1)}$, $x,y\leq N$ and then simply comparing coefficients. Alternatively, one can iterate work of Granville on bounds of coefficients of factors of single variable polynomials [Gra90]. ∎
@@
where the last line restricts to residue classes for which $a_{i}n+b_{i}$ are pairwise co-prime for all $i$, and where we note that such residue classes exist by the assumption that $P$ takes squarefree values infinitely-often. To be more precise, let $p\mid q$ and note that there necessarily exists $i\neq j$ such that $\ell_{i}(n),\ell_{j}(n)\equiv 0\ (\text{mod}\ p),$ for some $n$. Moreover, if both $i,j$ are such that $(a_{i},p)=1=(a_{j},p)$, then $n\equiv-b_{i}a_{i}^{-1},-b_{j}a_{j}^{-1}\ (\text{mod}\ p)$. If $(a_{i},p)=p$, say, then necessarily $p\mid b_{i}$ and $(a_{j},p)=1$ (otherwise, we would have that $p\mid b_{j}$, which is a contradiction to the fact that $P$ takes squarefree values infinitely-often); in particular, $n\equiv-b_{j}a_{j}^{-1}\ (\text{mod}\ p)$ and $p\mid\ell_{i}(n)$ for all $n$. Thus, we restrict to $m\ (\text{mod}\ q)$ which are not lifts to solutions of the above form, which yields
@@ 3. Products of Polynomials and Searching for Squares
(3.2) $\sum_{\begin{subarray}{c}{n_{1},n_{2},n_{3},n_{4}\leq N}\\ {d_{1},d_{2}\leq D}\end{subarray}}\operatorname{\mathds{1}}_{P(n_{1})P(n_{2})P(n_{3})P(n_{4})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}}+\sum_{\begin{subarray}{c}{n_{1},n_{2},n_{3},n_{4}\leq N}\\ {\min\{d_{1},d_{2}\}>D}\end{subarray}}\operatorname{\mathds{1}}_{P(n_{1})P(n_{2})P(n_{3})P(n_{4})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}}.$
@@ 3.3. Picking Parameters
$=\frac{1}{\left(\operatorname{\mathbb{E}}\left[M_{N}^{2}\right]\right)^{2}}\sum_{\begin{subarray}{c}{p,q\leq k_{N}}\\ {p\neq q}\end{subarray}}\left(\sum_{\begin{subarray}{c}{m_{1},m_{2},n_{1},n_{2}\leq N}\\ {P^{+}(P(m_{i}))=p,P^{+}(P(n_{i}))=q}\end{subarray}}\operatorname{\mathds{1}}_{P(m_{1})P(m_{2})P(n_{1})P(n_{2})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}}+O_{P}^{(p,q)}(1)\right)$
$\lesssim\frac{1}{(\kappa_{P}N)^{2}}\sum_{\begin{subarray}{c}{p,q\leq k_{N}}\\ {p\neq q}\end{subarray}}\sum_{\begin{subarray}{c}{m_{1},m_{2},n_{1},n_{2}\leq N}\\ {P^{+}(P(m_{i}))=p,P^{+}(P(n_{i}))=q}\end{subarray}}\operatorname{\mathds{1}}_{P(m_{1})P(m_{2})P(n_{1})P(n_{2})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}}$
$\lesssim 1+\frac{1}{(\kappa_{P}N)^{2}}\left(\sum_{\begin{subarray}{c}{p,q\leq k_{N}}\\ {p\neq q}\end{subarray}}\sum_{\begin{subarray}{c}{m_{1},m_{2},n_{1},n_{2}\leq N}\\ {P^{+}(P(m_{i}))=p,P^{+}(P(n_{i}))=q}\\ {m_{1}\neq m_{2},n_{1}\neq n_{2}}\end{subarray}}\operatorname{\mathds{1}}_{P(m_{1})P(m_{2})P(n_{1})P(n_{2})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}}\right),$
where $O_{P}^{(p,q)}(1)$ indicates the number of non-injective values of $P(m),P(n)$ whose largest prime factor is $p$ and $q$, respectively and where $\lesssim$ indicates upper bounds which hold up to $o_{N\rightarrow\infty}(1)$. Thus, it suffices to show that
@@
For the smooth sum, we use the fact that there are very few $p$-smooth numbers of size $N^{\deg P}$ with $p\leq\log\log N$. Indeed, if $\psi(x,y)$ denotes the number of $y$-smooth integers up to $x$, then we know that $\psi(x,y)=xu^{-(1+o(1))u}$ uniformly in $u\leq y^{1-\epsilon}$, for any $\epsilon>0,$ where $u=\log x/\log y;$ see [HT93, Corollary 1.3], for example. For the second sum, we use the second divisor bound from Lemma 2.2, which yields $\ll N^{2}/p^{2}+N/p+1$; summing over the primes $\log\log N\leq p\leq k_{N}$ then gives an admissible error (recalling that $k_{N}\ll_{P}N$ when $P$ is a product of linear factors, with $k_{N}\ll_{P}N^{2}$ if $P$ is an irreducible quadratic). Thus,
@@ 5.1. Proof Strategy
Conditioning on the primes $\ll N/\log N$, the above sum which, we denote by $S(N)$, may be viewed as $S(N)\approx\sum_{p\gg N\log N}f(p)c_{p}(N)$, where $c_{P}(N)$ are “fixed” coefficients. This is a weighted sum of independent and identically distributed random variables, which should have approximately Gaussian behavior, as long as the weights are not too irregular.
@@
Let $m\in\operatorname{\mathbb{N}}$ and let $S\subset\operatorname{\mathbb{N}}$ be finite and nonempty. Suppose that for each $1\leq i\leq m$ and $h\in S$ we are given a deterministic coefficient $c(i,h)\in\operatorname{\mathbb{C}}$. Finally, suppose that $(V_{i})_{1\leq i\leq m}$ is a sequence of independent, mean zero, complex-valued random variables, and let $Y=(Y_{h})_{h\in S}$ be the #$\#S$-dimensional random vector with components $Y_{h}:=\Re(\sum_{i=1}^{m}c(i,h)V_{i})$. If $Z=(Z_{h})_{h\in S}$ is a multivariate normal random vector with the same mean vector and covariance matrix as $Y$, then for any $u\in\operatorname{\mathbb{R}}$ and any small $\eta>0$, we have
$\mathbb{P}\left(\max_{h\in S}Y_{h}\leq u\right)$ $\leq\mathbb{P}\left(\max_{h\in S}Z_{h}\leq u+\eta\right)$
$\quad+O\left(\frac{1}{\eta^{2}}\sum_{g,h\in S}\sqrt{\sum_{i=1}^{m}|c(i,g)|^{2}|c(i,h)|^{2}\operatorname{\mathbb{E}}\left[|V_{i}|^{4}\right]}+\frac{1}{\eta^{3}}\sum_{i=1}^{m}\operatorname{\mathbb{E}}\left[|V_{i}|^{3}\right]\left(\sum_{h\in S}|c(i,h)|\right)^{3}\right).$
Using the notation of the lemma, we will choose: $V_{i}$ to be, essentially, the $f(p)$ for $p\gg N\log N$; $S$ to be the set of the indices of the $\lfloor 0.99k\rfloor$ scales; $c(i,h)$ to be the coefficient $c_{p}(x_{h})$ from the sentence following (5.1) (where $i$ corresponds to $p$). Note that both the set $S$ and the coefficients $c(i,h)$ are indeed deterministic upon conditioning.
@@
Suppose that $n\geq 2$, and that $\varepsilon\geq 0$ is sufficiently small (i.e., less than a certain small absolute constant). Let $X_{1},\dots,X_{k}$ be mean zero, variance one, jointly normal random variables, and suppose $\operatorname{\mathbb{E}}[X_{i}X_{j}]\leq\varepsilon$ whenever $i\neq j$. Then, for any $100100\varepsilon\leq\delta\leq 1/$ 100 (say), we have

### v2

@@ Abstract.
We study the distribution of partial sums of Rademacher random multiplicative functions $(f(n))_{n}$ evaluated at polynomial arguments. We show that for a polynomial $P\in\operatorname{\mathbb{Z}}[x]$ that is a product of at least two distinct linear factors or an irreducible quadratic satisfying a natural condition, there exists a constant $\kappa_{P}>0$ such that
@@ 1.1. Random Multiplicative Functions
Note that $f(n)$ is a random model for $\mu(n),$ as $\mu(n)$ is simply a specific realization of $f(n)$. Using the theory of Dirichlet series, Wintner [Win44] showed that $\sum_{n}f(n)n^{-s}$ is almost always convergent for $\Re(s)>1/2.$ He further showed that, for all $\varepsilon>0$, both $\sum_{n\leq x}f(n)=O_{\varepsilon}(x^{1/2+\varepsilon})$ and $\sum_{n\leq x}f(n)\neq O_{\varepsilon}(x^{1/2-\epsilon})$ hold almost always, leading to the egregious statement that “the Riemann Hypothesis is almost surely true”.
@@ 1.2. The Random Chowla Conjecture
Let $f$ be a Rademacher RMF and let $P\in\operatorname{\mathbb{Z}}[x]$ be a product of at least two distinct linear factors over $\operatorname{\mathbb{Z}}$ or irreducible of degree $2$. Suppose that there is no prime $p$ for which $p^{2}\mid P(n)$ for all $n\in\operatorname{\mathbb{N}}$. Then, there exists a constant $\kappa_{P}>0$ such that
@@
Remark 1.1.
If there is no prime $p$ such that $p^{2}\mid P(n)$ for all $n\in\operatorname{\mathbb{N}}$, then we call the polynomial $P$ admissible.
@@
respectively. The former is still “trivial”, as a product of two squarefree numbers is a perfect square if and only if the two numbers are equal; and so, the second moment amounts to counting the number of $n\leq N$ for which $P(n)$ is squarefree. This is a notoriously difficult problem for general polynomials, which we briefly discuss at the end of Section 2. The main technical result in this paper is proving an asymptotic for (1.3) (Proposition 3.1). The key new ingredient to our proof is a bootstrapping argument, where it is essential that our polynomials are either products of linear factors or irreducible quadratics.
@@ 1.4. Completely Multiplcative RMFs and Quadratic Twists
which counts the number of integral points up to a given height on the family of quadratic twists of the curve with affine model $C:y^{2}=P(x)$. In the case where $P$ has degree $2$, we are able to solve this counting problem using work of Hooley on Pell equations [Hoo84]: Theorem 1 of [Hoo84] tells us that almost all twists will have at most one integral point of low height, so that we can replace the second moment for these $d$ by the first moment, which exactly counts diagonal solutions $n_{1}=n_{2}$. For the remaining twists, trivial bounds suffice. For higher degree polynomials, analogues of Hooley’ results exist, conditionally on the $abc$-conjecture and for suitable $P$, by work of Granville [Gra07, Theorem 4 (i)]; as such, we are able to solve special cases of the second moment problem for completely multiplicative Rademacher RMFs. The fourth moment is substantially more complicated. For starters, one would like to know that almost all twists of the surface $y^{2}=P(x_{1})P(x_{2})$ have at most one integral point. Nevertheless, we still expect the behaviour to be Gaussian and this is a subject of ongoing investigation.
@@ 2.1. Martingales and McLeish’s CLT
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $(\mathcal{F}_{n})_{n}$ be a filtration on $(\Omega,\mathcal{F},\mathbb{P})$ (i.e., a sequence of sub-$\sigma$-algebras satisfying $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\dots\subseteq\mathcal{F}$). A sequence of random variables $(X_{n})_{n}$ is called a martingale difference sequence on $(\Omega,\mathcal{F},(\mathcal{F}_{n})_{n},\mathbb{P})$ if
@@ 2.2. Squarefree Values and Divisor Bounds
Let $Q\in\operatorname{\mathbb{Z}}[x]$, $y,N\in\operatorname{\mathbb{N}}$, and $\omega(y):=\#\{p:p\mid y\}$. If $y$ is squarefree and $Q$ is not the zero polynomial modulo $p$ for any prime $p\mid y$, then:
@@
Given $F(x,y),$ consider its prime factorization $f_{1}(x,y)\cdots f_{k}(x,y)$ in $\operatorname{\mathbb{C}}[x,y]$ and note that $F(x,y)=0$ if and only if $f_{i}(x,y)=0$ for some $i$; in particular, the number of integral points on $F(x,y)=0$ is the sum of the number of integral points on each $f_{i}(x,y)$. We note that none of the $f_{i}$’s have degree $1$; see e.g., [WX24, Lemma 2.5].
For $f_{i}$ with $\deg f_{i}\geq 3$, there are at most $\ll_{\varepsilon}N^{1/\deg f_{i}+\varepsilon}\ll N^{1/3+\varepsilon}$ integral points in the desired box, by Lemma 2.3.
For $f_{i}$ of degree $2$, we may assume, without loss of generality, that $f_{i}(x,y)\in\operatorname{\mathbb{Q}}[x,y]$: if not, we note that $f_{i}$ comes with a conjugate pair [^3] Since $F(x,y)\in\operatorname{\mathbb{Z}}[x,y],$ if $f_{i}$ has degree $2$ and $f_{i}\not\in\operatorname{\mathbb{Q}}[x,y]$, then there must be an $f_{j}$ such that $f_{i}f_{j}\in\operatorname{\mathbb{Q}}[x,y].$, say $f_{j}$, as $F(x,y)\in\operatorname{\mathbb{Z}}[x,y];$ in particular, the number of rational points on $f_{i}(x,y)=0$ is equal to the number of rational points on the intersection of the curves $f_{i}(x,y)=0=f_{j}(x,y)$, which is at most $4$ by Bezout’s Theorem on intersections of plane curves [Béz79]. In the case where $F(x,y)$ has quadratic factors in $\operatorname{\mathbb{Q}}[x,y]$, then the coefficients of this quadratic factor are $\ll N^{O_{\deg P}(1)}$ whenever $a,b\ll N^{O_{\deg P}(1)}$. To see this, note that the left hand side of $F(x,y)=f_{1}(x,y)\cdots f_{k}(x,y)$ is at most $N^{O_{\deg P}(1)}$ for $a,b\ll N^{O_{\deg P}(1)}$, $x,y\leq N$; comparing coefficients yields the desired result. Alternatively, one can iterate work of Granville on bounds of coefficients of factors of single variable polynomials [Gra90]. If the discriminant of the quadratic factor is not a square, we apply Lemma 2.4 and get that there are at most $\ll_{\varepsilon}N^{\varepsilon}$ integral points across such quadratic irreducible factors. In the case where it is a square, then the quadratic terms can be written as a product of two linear factors plus a constant term, say $\ell_{1}(x,y)\ell_{2}(x,y)+g$, with $g$ necessarily non-zero; and so, there are at most $\ll_{\varepsilon}N^{\varepsilon}$ integral points on $\ell_{1}(x,y)\ell_{2}(x,y)+g=0$ by the divisor bound and we are done. ∎
@@
where we have used the fact that for all $m\ (\text{mod}\ q)$ and all $i\neq j$, either $(\ell_{i}(n),\ell_{j}(n))=1$ for all $n\equiv m\ (\text{mod}\ q)$ or $(\ell_{i}(n),\ell_{j}(n))>1$ for all $n\equiv m\ (\text{mod}\ q)$. To see this, note that if $m$ and $n$ are such that $(\ell_{i}(n),\ell_{j}(n))>1$ with $n\equiv m\ (\text{mod}\ q)$, then there exists a prime $p$ such that $p\nmid q$, by definition of $q$. Furthermore, $n\equiv-b_{i}a_{i}^{-1}\ (\text{mod}\ p)$, say, as at least one of $a_{i},a_{j}$ must not be divisible by $p$, by the assumption that $P$ is admissible. Hence, for any other $n_{0}\equiv m\ (\text{mod}\ q)$, we have that $n_{0}\equiv m\ (\text{mod}\ p)$ with $m\equiv-b_{i}a_{i}^{-1}\ (\text{mod}\ p)$, so that $(\ell_{i}(n_{0}),\ell_{j}(n_{0}))>1$ and we are done. In particular,
@@ 3. Products of Polynomials and Searching for Squares
(3.2) $\sum_{\begin{subarray}{c}{n_{1},n_{2},n_{3},n_{4}\leq N}\\ {d_{1},d_{2}\leq D}\end{subarray}}\operatorname{\mathds{1}}_{P(n_{1})P(n_{2})P(n_{3})P(n_{4})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}}+\sum_{\begin{subarray}{c}{n_{1},n_{2},n_{3},n_{4}\leq N}\\ {\min\{d_{1},d_{2}\}>D}\end{subarray}}\operatorname{\mathds{1}}_{P(n_{1})P(n_{2})P(n_{3})P(n_{4})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}},$
where emphasize that the first sum is over all $n_{1},n_{2},n_{3},n_{4}\leq N$ for which $d_{1}=(P(n_{1}),P(n_{2}))$ and $d_{2}=(P(n_{3}),P(n_{4}))$ are $\leq D$ and similarly for the second sum.
@@ 3.3. Picking Parameters
$=\frac{1}{\left(\operatorname{\mathbb{E}}\left[M_{N}^{2}\right]\right)^{2}}\sum_{\begin{subarray}{c}{p,q\leq k_{N}}\\ {p\neq q}\end{subarray}}\left(\sum_{\begin{subarray}{c}{m_{1},m_{2},n_{1},n_{2}\leq N}\\ {P^{+}(P(m_{i}))=p,P^{+}(P(n_{i}))=q}\end{subarray}}\operatorname{\mathds{1}}_{P(m_{1})P(m_{2})P(n_{1})P(n_{2})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}}\right)$
$=1+o(1)+\frac{1}{\left(\operatorname{\mathbb{E}}\left[M_{N}^{2}\right]\right)^{2}}\left(\sum_{\begin{subarray}{c}{p,q\leq k_{N}}\\ {p\neq q}\end{subarray}}\sum_{\begin{subarray}{c}{m_{1},m_{2},n_{1},n_{2}\leq N}\\ {P^{+}(P(m_{i}))=p,P^{+}(P(n_{i}))=q}\\ {m_{1}\neq m_{2},n_{1}\neq n_{2}}\end{subarray}}\operatorname{\mathds{1}}_{P(m_{1})P(m_{2})P(n_{1})P(n_{2})={\raisebox{-1.0pt}{$\square$}}}^{\text{SF}}\right),$
where the $o(1)$ term takes into account the non-injective values of $P$. Since $\operatorname{\mathbb{E}}\left[M_{N}^{2}\right]\sim\kappa_{P}N$, it suffices to show that
@@
For the smooth sum, we use the fact that there are very few $p$-smooth numbers of size $N^{\deg P}$ with $p\leq\log\log N$. Indeed, if $\psi(x,y)$ denotes the number of $y$-smooth integers up to $x$, then we know that $\psi(x,y)=xu^{-(1+o(1))u}$ uniformly in $u\leq y^{1-\epsilon}$, for any $\epsilon>0,$ where $u=\log x/\log y;$ see [HT93, Corollary 1.3], for example. This results in an error of size $\ll_{\varepsilon}N^{\varepsilon}$. For the second sum, we split into cases, depending on the size of $p$. For the primes $p\ll_{P}N$ we use the second divisor bound from Lemma 2.2 which yields the upper bound $\ll\left(N^{2}/p^{2}+N/p+1\right)^{2}$ on the inner sum. Summing over the small primes then yields an admissible error. For the larger primes, note that there are $\ll_{P}N$ primes $p\gg N$ such that $p\mid P(n)$ for some $n\leq N$ (since each such value of $P(n)$ is divisible by $\ll\deg P$ of such large primes). For each of these primes, the number of $n\leq N$ such that $P(n)\equiv 0\pmod{p}$ is $\ll 1$, again by Lemma 2.2, which yields a total error of $\ll_{P}N$. Thus,
@@ 5.1. Proof Strategy
Conditioning on the primes $\ll N/\log N$, the above sum which, we denote by $S(N)$, may be viewed as $S(N)\approx\sum_{p\gg N\log N}f(p)c_{p}(N)$, where $c_{P}(N):=\sum_{n\leq N:p\mid n^{2}+1}f\left(\frac{n^{2}+1}{p}\right)$ are “fixed” coefficients. This is a weighted sum of independent and identically distributed random variables, which should have approximately Gaussian behavior, as long as the weights are not too irregular.
@@
Let $m\in\operatorname{\mathbb{N}}$ and let $S\subset\operatorname{\mathbb{N}}$ be finite and nonempty. Suppose that for each $1\leq\ell\leq m$ and $h\in S$ we are given a deterministic coefficient $c(\ell,i)\in\operatorname{\mathbb{C}}$. Finally, suppose that $(V_{\ell})_{1\leq\ell\leq m}$ is a sequence of independent, mean zero, complex-valued random variables, and let $Y=(Y_{i})_{i\in S}$ be the #$\#S$-dimensional random vector with components $Y_{i}:=\Re(\sum_{\ell=1}^{m}c(\ell,i)V_{\ell})$. If $Z=(Z_{i})_{i\in S}$ is a multivariate normal random vector with the same mean vector and covariance matrix as $Y$, then for any $u\in\operatorname{\mathbb{R}}$ and any small $\eta>0$, we have
$\mathbb{P}\left(\max_{i\in S}Y_{i}\leq u\right)$ $\leq\mathbb{P}\left(\max_{i\in S}Z_{i}\leq u+\eta\right)$
$\quad+O\left(\frac{1}{\eta^{2}}\sum_{i,j\in S}\sqrt{\sum_{\ell=1}^{m}|c(\ell,i)|^{2}|c(\ell,j)|^{2}\operatorname{\mathbb{E}}\left[|V_{\ell}|^{4}\right]}+\frac{1}{\eta^{3}}\sum_{\ell=1}^{m}\operatorname{\mathbb{E}}\left[|V_{\ell}|^{3}\right]\left(\sum_{i\in S}|c(\ell,i)|\right)^{3}\right).$
Using the notation of the lemma, we will choose: $V_{\ell}$ to be, essentially, the $f(p)$ for $p\gg N\log N$; $S$ to be the set of the indices of the $\lfloor 0.99k\rfloor$ scales; $c(\ell,i)$ to be the coefficient $c_{p}(x_{i})$ from the sentence following (5.1) (where $\ell$ corresponds to $p$). Note that both the set $S$ and the coefficients $c(\ell,i)$ are indeed deterministic upon conditioning.
@@
Suppose that $n\geq 2$, and that $\varepsilon\geq 0$ is sufficiently small (i.e., less than a certain small absolute constant). Let $X_{1},\dots,X_{k}$ be mean zero, variance one, jointly normal random variables [^8] We say that two random variables $X,Y$ are jointly normal if $aX+bY$ is a normal random variable for all $a,b\in\operatorname{\mathbb{R}}$., and suppose $\operatorname{\mathbb{E}}[X_{i}X_{j}]\leq\varepsilon$ whenever $i\neq j$. Then, for any $100100\varepsilon\leq\delta\leq 1/$ 100 (say), we have